OpenStudy (anonymous):

Factor completely 3x4 – 6x3 + 12x2, Explain?

OpenStudy (whpalmer4):

First, factor out any common factors, either numeric or in the variable.

OpenStudy (whpalmer4):

What do you get if you do that?

OpenStudy (anonymous):

3x^2(x^2-2x+4) ?

OpenStudy (whpalmer4):

\[3x^4-6x3 +12x^2 = 3x^2(x^2-2x+4)\] Now, can you factor the polynomial in the parentheses?

OpenStudy (anonymous):

yes?

OpenStudy (whpalmer4):

How does it factor?

OpenStudy (anonymous):

3x2(x – 2)(x – 2) ?

OpenStudy (whpalmer4):

hmm. what do you get if you multiply \[(x-2)(x-2)\]

OpenStudy (anonymous):

3x^2(x^2-4x+4) so it cant be facotr further?

OpenStudy (whpalmer4):

In fact, this cannot be factored further, as \(x^2-2x+4\) is irreducible. So, \[3x^4-6x^3+12x^2 = 3x^2(x^2-2x+4)\]is a complete factoring.

OpenStudy (whpalmer4):

If we put \(x^2-2x+4=0\) through the quadratic formula, we see the solutions (which are the same as the factors, of course) are \[x=\frac{-(-2)\pm\sqrt{(-2)^2-4(1)(4)}}{2(1)} =\frac{4\pm\sqrt{4-16 }}{2} = 2\pm\sqrt{3} \]

OpenStudy (whpalmer4):

so it could be factored as \[3x^2(x-(2+\sqrt{3}))(x-(2-\sqrt{3}))\]but typically that isn't done.

OpenStudy (whpalmer4):

ah, crap, I forgot the \(i\)s! :-( \[x = 2 \pm i\sqrt{3}\] \[3x^2(x-(2+i\sqrt{3}))(x-(2-i\sqrt{3}))\] Thought that was a little too easy to type :-)

OpenStudy (anonymous):

thanks